40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Mon Aug 16 19:12:18 2010, Jenne, Update, SUS, More TT characterization 
    Reply  Mon Aug 16 23:39:29 2010, Yoichi, Update, SUS, More TT characterization EQStopScrewHoles.jpg
Message ID: 3425     Entry time: Mon Aug 16 19:12:18 2010     Reply to this: 3427
Author: Jenne 
Type: Update 
Category: SUS 
Subject: More TT characterization 

[Jenne, Yoichi]

We characterized Tip Tilts numbers 2 & 3 today.  Recall #4 is the one which Koji and I measured some time ago, and #s 1 & 5 have yet to be suspended (that's on the to-do list for tomorrow).

When we began looking at #3 (the one which had been in the BS chamber for a few days, but was removed for characterization) we found that the pitch pointing was way off.  The beam was way too low after reflection.  So we adjusted that (and got it right on the first try....a miracle!).  This does however make me pretty concerned about our in-chamber pointing.  Are we destroying our pointing during travel between the cleanroom and the chambers?  Is there anything we can do about it? Pointing doesn't seem to be lost when we move them around on the tables in the cleanroom, ie we can pick up a TT, move it, leave it for a while, move it back to the oplev, and the pointing still seems okay.  But the TT which was sent to the chambers came back with bad pointing. I'm sure this is related to the hysteresis we see in the pointing if we poke the top of the mirror holder versus the bottom when exciting pitch motion.

For both #2 and #3, we measured the frequency and Q of Pitch, Yaw, Pos, Side, Vert motion.  For the Vert motion, we measured both without and with EQ stops as dampers.  For the other modes, all were measured with the ECD backplane in place.  Pitch and Yaw were measured with reflection off of the mirror surface onto the PD, while Pos, Side and Vert were measured using the wire clamp on the mirror holder to obscure the beam as a shadow sensor.

TT #2

Pitch: Overdamped, no freq measured, Q < 1/2

Yaw: freq ~1.8Hz, Q between 2-7

Pos: freq ~1.75Hz, Q too low to measure, but above critically damped

Side: freq ~ 1.8Hz, Q~5

Vert no dampers: freq ~20Hz, Q~36

Vert with dampers on outer screws: freq~24Hz, Q~8,

TT #3

Pitch: no freq measured.  Q~1/2?  Upon being excited in Pitch, the beam started down way below the photodiode, came up a little past its DC place, and went back down a tiny bit.  So not quite overdamped.

Yaw: freq ~1.96Hz, Q very low

Pos: freq ~1.72Hz, Q~3

Side: freq ~1.85Hz, Q~6

Vert no dampers: freq ~20Hz, Q~75

Vert with dampers on outer screws: freq ~20Hz, Q~34  (Frequency stayed constant....we did several measurements both with and without the dampers...but the half life time changed significantly)

 

Things we noticed:  Koji and I had been concerned the last time we were looking at TT#2 because the frequency got lower and the Q seemed to get higher when we added the damping to the vertical blades.  Yoichi and I did not find that to be true today.  We did notice, however, that the EQ stop screws with the viton had been placed in the holes closer to the clamping point, whereas with TT #4 the screws had been placed in the holes farther from the clamping point.  We moved the screws on TT #2 to the outer holes, and noticed that the frequency increased slightly, and the Q significantly decreased.  We then followed this outer-hole philosophy when installing screws in TT #3.

To Do List: We need to suspend the ECDs and the Optics for the remaining two Tip Tilts, and to characterize them.  We also (probably farther-future) need to take transfer functions using a shaker / shake table with our spare Tip Tilt.  After all the TTs are suspended and have their modes measured, we will be ready for installation into the chamber during the next vent.

 

ELOG V3.1.3-